OpenStudy (anonymous):

Each of the letters in the word MATHEMATICS is represented by a letter tile in a bag. Stacey picks three letters, without replacement. To the nearest whole percent, which is the probability that she'll get all vowels? ANSWER CHOICES 2%-3%-5%-6%

OpenStudy (anonymous):

how many letters in "mathematics" ?

OpenStudy (anonymous):

guess i have to count for myself damn

OpenStudy (anonymous):

11 lol

OpenStudy (anonymous):

i get eleven four vowels probability first one is a vowel is \(\frac{4}{11}\) probability second one is a vowel given that the first on is \(\frac{3}{10}\) probability that third is given etc is \(\frac{2}{9}\) and last one is \(\frac{1}{8}\) multiply all these for your answer

OpenStudy (anonymous):

\[\frac{ 4 }{ 11 }\times \frac{ 3 }{ 10 }\times \frac{ 2 }{ 9 }=\frac{ 4 }{165 }\]

OpenStudy (anonymous):

oops pick 3 letters go with @pgpilot326 answer, not mine

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome!